# LiDMaS: Architecture-Level Modeling of Fault-Tolerant Magic-State Injection in GKP Photonic Qubits

### v1

@@ III.5 Phase-boundary analysis
Finally, we construct phase-boundary plots identifying the minimum squeezing required to simultaneously satisfy a target success probability and logical fidelity. Figure 7 shows the boundary defined by the conditions $P_{\mathrm{success}}\geq 0.95$,   $F_{\mathrm{logical}}\geq 0.79$, for different surface-code distances.
The phase boundary reveals a clear trade-off between squeezing resources and outer-code distance. Increasing the code distance systematically shifts the boundary toward lower squeezing, demonstrating that stronger logical protection can partially compensate for limited squeezing resources. This representation provides a compact and experimentally relevant design chart for photonic fault-tolerant architectures, enabling direct comparison of hardware capabilities and error-correction requirements.
Figure 7: Phase boundary showing the minimum squeezing required to achieve both high RUS success probability ($\geq 0.95$) and high logical fidelity ($\geq 0.79$) as a function of baseline loss. Higher surface-code distances reduce the squeezing requirement, illustrating the trade-off between hardware quality and error-correction overhead.
@@ IV Discussions
A major line of work on fault-tolerant photonic quantum computation has focused on the Gottesman–Kitaev–Preskill (GKP) encoding under finite squeezing, most notably in the seminal analysis by Menicucci et al. [14], which established threshold theorems for continuous-variable cluster-state computation using idealized Gaussian operations. That study rigorously analyzed error propagation in measurement-based quantum computing and demonstrated that fault tolerance is achievable in principle with sufficiently high squeezing. However, the analysis primarily emphasized asymptotic thresholds and did not explicitly consider magic-state injection protocols, repeat-until-success dynamics, or the impact of heralded photon loss on gate-level overhead. In contrast, our work directly models logical $T$-gate magic-state preparation, explicitly incorporating finite squeezing, depolarizing noise, and erasure into a unified architecture-level framework. Rather than focusing on global thresholds, we quantify operational metrics, success probability, injection overhead, and logical fidelity, relevant to near-term photonic architectures.
@@
Recent studies by Fukui et al. [8] and related follow-up work introduced high-performance GKP error-correction schemes based on analog information, demonstrating that exploiting continuous measurement outcomes can substantially improve logical error suppression at moderate squeezing levels. These works provide important insights into decoder-level performance and show that finite squeezing noise can be mitigated more efficiently than previously thought. However, they largely abstract away from higher-level logical workflows such as magic-state distillation or injection, and they do not address how residual errors interact with outer stabilizer codes under realistic loss conditions. Our approach complements these results by treating GKP protection as an effective logical noise map and explicitly studying how residual errors propagate through a non-Clifford injection protocol, thereby bridging the gap between decoder-centric analyses and system-level fault-tolerant gate synthesis.
@@
Magic-state preparation in photonic systems has also been explored in detail by Campbell, Browne [5], and collaborators, particularly in the context of measurement-based implementations and resource-efficient non-Clifford gate synthesis. These studies emphasize the theoretical optimality of magic-state distillation protocols and their overhead scaling in qubit-based architectures. While these results are foundational, they typically assume abstract error models and do not explicitly incorporate photonic constraints such as finite squeezing or photon-loss-induced erasure. In contrast, our work focuses on injection rather than distillation and shows that, in photonic architectures, high logical fidelity can be achieved with minimal overhead through RUS-style injection combined with outer code protection. This highlights an alternative pathway in which distillation-heavy pipelines may be deferred or partially replaced by high-quality injected states in sufficiently well-engineered photonic platforms.
@@
Surface-code-protected photonic architectures have been proposed by Nickerson, Li, and Benjamin [15], who analyzed fault-tolerant quantum computation in modular and networked systems subject to photon loss. Their work demonstrated that heralded loss can be tolerated efficiently using erasure-aware decoding, significantly relaxing hardware requirements. However, their analysis focused primarily on Clifford operations and logical error rates, without explicitly examining non-Clifford gate synthesis or magic-state fidelity. Our results align with their conclusions regarding loss tolerance but extend them by showing that, once heralded loss is accounted for, logical magic-state fidelity becomes largely insensitive to moderate loss levels. Instead, finite squeezing emerges as the dominant factor limiting logical quality, a distinction that is made explicit through our sensitivity and phase-boundary analyses.
@@
Finally, recent photonic fault-tolerance roadmaps, such as those developed by Rudolph [19] and subsequent architectural reviews, have emphasized the central role of squeezing as a hardware bottleneck while acknowledging uncertainty in how squeezing, loss, and code distance jointly constrain performance. These works provide broad qualitative guidance but stop short of quantitative design rules. Our phase-boundary diagrams directly address this gap by identifying the minimum squeezing required to simultaneously achieve high RUS success probability and high logical fidelity across different loss regimes and code distances. By mapping out these boundaries explicitly, our study transforms qualitative architectural intuition into quantitative constraints, offering practical guidance for experimental photonic quantum computing platforms.
@@ V Conclusion
In this work, we have presented an architecture-level study of fault-tolerant logical magic-state preparation in GKP-encoded photonic quantum computing architectures. By combining a repeat-until-success (RUS) logical $T$-gate injection protocol with an abstracted outer surface-code layer, we systematically quantified the interplay between finite squeezing, photon-loss-induced erasure, and code distance. Our approach enables the efficient exploration of a broad design space that would be prohibitively expensive to access using full continuous-variable wavefunction simulations or decoder-level surface-code modeling.
Across all regimes studied, we find that logical magic-state injection remains robust against moderate photon loss, with the average RUS overhead remaining close to unity even in relatively lossy settings. In contrast, finite squeezing emerges as the dominant continuous error source affecting logical fidelity. Once protected by an outer code, logical magic-state fidelity rapidly saturates with increasing code distance, exhibiting only weak dependence on loss. Sensitivity analysis confirms this separation of roles, demonstrating that squeezing primarily governs logical quality, while loss predominantly impacts heralded success probabilities rather than post-selected fidelity.
Phase-boundary analysis further reveals clear threshold-like behavior, allowing us to identify the minimum squeezing requirements necessary to simultaneously achieve high success probability and high logical fidelity as a function of loss and code distance. These boundaries provide concrete quantitative guidance for the co-design of photonic hardware and fault-tolerant logical architectures.
All simulations in this work were performed using a custom density-matrix-based simulator implemented directly in Python and NumPy. No external quantum simulation frameworks, such as PennyLane, were employed. This design choice was deliberate: the objective of this study is to isolate architectural and logical trade-offs using analytically motivated noise abstractions, rather than to model device-level continuous-variable dynamics. The resulting framework is lightweight, transparent, and readily extensible, making it well suited for rapid architectural exploration.
Results clarify the relative importance of squeezing, loss, and code distance in photonic fault-tolerant quantum computing and provide a scalable modeling framework that complements more detailed device-level simulations. We anticipate that this approach will serve as a useful foundation for future studies bridging logical architecture design and experimental photonic platforms.
@@
No external quantum simulation frameworks or hardware backends were employed; instead, a custom density-matrix-based simulator was developed to model logical quantum operations, noise channels, and fault-tolerant protocols at the architectural level. This lightweight approach enabled efficient exploration of large parameter spaces relevant to photonic fault-tolerant quantum computing.
Any opinions, findings, conclusions, or recommendations expressed in this research are those of the author(s) and do not necessarily reflect the views of their respective affiliations.

### v3

@@ III.5 Phase-boundary analysis
Finally, phase-boundary plots are constructed to identify the minimum squeezing required to simultaneously satisfy a target success probability and logical fidelity. Figure 8 shows the boundary defined by the conditions $P_{\mathrm{success}}\geq 0.95$,   $F_{\mathrm{logical}}\geq 0.79$, for different surface-code distances.
The phase boundary reveals a clear trade-off between squeezing resources and outer-code distance. Increasing the code distance systematically shifts the boundary toward lower squeezing, demonstrating that stronger logical protection can partially compensate for limited squeezing resources; for $d=1$ no points satisfy both thresholds within the simulated range. This representation provides a compact and experimentally relevant design chart for photonic fault-tolerant architectures, enabling direct comparison of hardware capabilities and error-correction requirements.
Figure 8: Phase boundary showing the minimum squeezing required to achieve both high RUS success probability ($\geq 0.95$) and high logical fidelity ($\geq 0.79$) as a function of baseline loss. Higher surface-code distances reduce the squeezing requirement, illustrating the trade-off between hardware quality and error-correction overhead; for $d=1$ no feasible points meet both thresholds in the simulated range.
@@ IV Discussions
A major line of work on fault-tolerant photonic quantum computation has focused on the Gottesman–Kitaev–Preskill (GKP) encoding under finite squeezing, most notably in the seminal analysis by Menicucci et al. [14], which established threshold theorems for continuous-variable cluster-state computation using idealized Gaussian operations. That study rigorously analyzed error propagation in measurement-based quantum computing and demonstrated that fault tolerance is achievable in principle with sufficiently high squeezing. However, the analysis primarily emphasized asymptotic thresholds and did not explicitly consider magic-state injection protocols, repeat-until-success dynamics, or the impact of heralded photon loss on gate-level overhead (Table 1). In contrast, our work directly models logical $T$-gate magic-state preparation, explicitly incorporating finite squeezing, depolarizing noise, and erasure into a unified architecture-level framework. Rather than focusing on global thresholds, operational metrics—success probability, injection overhead, and logical fidelity—are quantified to provide performance measures relevant to near-term photonic architectures.
Table 1 summarizes how this work relates to representative prior studies on GKP fault tolerance, photonic architectures, and magic-state preparation.
Table 1: Comparison with representative literature on GKP fault tolerance, photonic architectures, and magic-state preparation.
Work Platform / code Primary focus Non-Clifford handling Loss treatment Metrics
Menicucci et al. (2014) [14] CV cluster states with GKP Fault-tolerance thresholds for MBQC Universal MBQC (no injection) Not central Thresholds vs squeezing
Fukui et al. (2018) [8] GKP with analog QEC Decoder-level thresholds Not explicit Not central Logical error rates
Noh & Chamberland (2020) [16] Surface–GKP code Bosonic FT QEC Not explicit Not central Logical error scaling
Nickerson et al. (2014) [15] Photonic modular architecture Loss-tolerant networking Not explicit Erasure-aware Loss thresholds, logical rates
Bourassa et al. (2021) [2] Photonic GKP architecture System-level blueprint Distillation at high level Loss-aware Resource/ overhead estimates
Campbell et al. (2012) [4] Qubit codes (Reed–Muller) Magic-state distillation Distillation protocols Not central Distillation overheads
This work GKP photonic + surface abstraction RUS-based injection architecture Explicit RUS $T$ injection Heralded erasure $P_{\mathrm{succ}}$, $\langle R\rangle_{\mathrm{succ}}$, $F_{\mathrm{log}}$
@@
Recent studies by Fukui et al. [8] and related follow-up work introduced high-performance GKP error-correction schemes based on analog information, demonstrating that exploiting continuous measurement outcomes can substantially improve logical error suppression at moderate squeezing levels. These works provide important insights into decoder-level performance and show that finite squeezing noise can be mitigated more efficiently than previously thought. Related bosonic fault-tolerance analyses using surface–GKP codes, such as Noh and Chamberland [16], report logical-error scaling but likewise do not model non-Clifford injection or loss in detail (Table 1). Our approach complements these results by treating GKP protection as an effective logical noise map and explicitly studying how residual errors propagate through a non-Clifford injection protocol, thereby bridging the gap between decoder-centric analyses and system-level fault-tolerant gate synthesis.
@@
Magic-state preparation has been studied extensively in qubit-based architectures, particularly through distillation protocols and measurement-based constructions [5, 3, 4, 13]. These studies emphasize theoretical optimality and overhead scaling under abstract noise models, and they do not explicitly incorporate photonic constraints such as finite squeezing or photon-loss-induced erasure (Table 1). In contrast, our work focuses on injection rather than distillation and shows that, in photonic architectures, high logical fidelity can be achieved with modest overhead through RUS-style injection combined with outer-code protection within the simulated regime. This suggests an alternative pathway in which distillation-heavy pipelines could be reduced when injected-state quality is sufficiently high.
@@
Surface-code-protected photonic architectures have been proposed by Nickerson, Li, and Benjamin [15], who analyzed fault-tolerant quantum computation in modular and networked systems subject to photon loss. Their work demonstrated that heralded loss can be tolerated efficiently using erasure-aware decoding, significantly relaxing hardware requirements (Table 1). However, their analysis focused primarily on Clifford operations and logical error rates, without explicitly examining non-Clifford gate synthesis or magic-state fidelity. Our results align with their conclusions regarding loss tolerance but extend them by showing that, once heralded loss is accounted for, logical magic-state fidelity is only weakly sensitive to moderate loss levels within the simulated range. Instead, finite squeezing emerges as the dominant factor limiting logical quality, a distinction that is made explicit through our sensitivity and phase-boundary analyses.
@@
Finally, recent photonic fault-tolerance roadmaps, such as those developed by Rudolph [19] and subsequent architectural reviews [2, 21], have emphasized the central role of squeezing as a hardware bottleneck while acknowledging uncertainty in how squeezing, loss, and code distance jointly constrain performance. These works provide broad qualitative guidance but stop short of quantitative design rules. Our phase-boundary diagrams address this gap by identifying the minimum squeezing required to simultaneously achieve high RUS success probability and high logical fidelity across different loss regimes and code distances. By mapping out these boundaries explicitly, our study helps translate qualitative architectural intuition into quantitative constraints, offering practical guidance for experimental photonic quantum computing platforms.
@@ V Conclusion
An architecture-level study of fault-tolerant logical magic-state preparation in GKP-encoded photonic quantum computing architectures has been presented. A repeat-until-success (RUS) logical $T$-gate injection protocol is combined with an abstracted outer surface-code layer to quantify how finite squeezing, photon-loss-induced erasure, and code distance jointly shape operational performance. This framework enables efficient exploration of a broad design space that would otherwise be prohibitively expensive to access using full continuous-variable wavefunction simulations or decoder-level surface-code modeling.
Within the simulated parameter ranges, logical magic-state injection is robust against moderate photon loss, with average RUS overheads remaining near unity even in relatively lossy settings. In contrast, finite squeezing is the dominant continuous error source affecting logical fidelity. Once protected by an outer code, logical magic-state fidelity shows only weak dependence on loss and diminishing gains with increasing code distance. Sensitivity analysis confirms this separation of roles: squeezing primarily governs logical quality, while loss predominantly impacts heralded success probabilities rather than post-selected fidelity.
Phase-boundary analysis reveals threshold-like boundaries in the simulated regime, identifying the minimum squeezing required to simultaneously achieve high success probability and high logical fidelity as functions of loss and code distance. These boundaries provide quantitative guidance for the co-design of photonic hardware and fault-tolerant logical architectures.
All simulations are performed using a custom density-matrix-based simulator implemented directly in Python and NumPy. No external quantum simulation frameworks, such as PennyLane, are employed. The framework is lightweight, transparent, and readily extensible, supporting rapid architectural exploration. These results clarify the relative importance of squeezing, loss, and code distance in photonic fault-tolerant quantum computing and provide a scalable modeling foundation that complements, rather than replaces, detailed device-level simulations.
@@
Any opinions, findings, conclusions, or recommendations expressed in this research are those of the author and do not necessarily reflect the views of their respective affiliations.